Salvatori, Inc., manufactures and sells two products: Product A4 and Product 05. Data concerning the expected production of eSalvatori, Inc., manufactures and sells two products: Product A4 and Product Q5. Data concerning the expected production of each product and the expected total direct labor-hours (DLHs) required to produce that output appear below: Expected Production Direct Labor-Hours Per Unit Total Direct Labor-Hours Product A4 650 7.5 4,875 Product Q5 950 4.5 4,275 Total direct labor-hours 9,150 The company has an activity-based costing system with the following activity cost pools, activity measures, and expected activity: Estimated Expected Activity Activity Cost Pools Activity Measures Overhead Cost Product A4 Product Q5 Total Labor-related DLHs $ 165,558 4,875 4,275 9,150 Machine setups setups 13,050 1,050 950 2,000 Order size MHs 502,027 4,700 5,000 9,700 $ 680,635 The overhead applied to each unit of Product A4 under activity-based costing is closest to: (Round your intermediate calculations to 2 decimal places.)
Answer #1

Answer is 520.49 approx.calculation is as below

Expected Prod Direct Labor hrs per unit Total DL hours
Product A4                        650.00                 7.50                    4,875.00
Product Q5                        950.00                 4.50                    4,275.00
                   1,600.00                    9,150.00
a b c d=b+c e=a/d f=e*b
Expected Activity
Activity Cost Pools Activity Measures Estimated OH cost Product A4 Product Q5 Total Estimated cost per activity Measure Estimated cost for Product A4
Labor Related DLHs    165,558.00                    4,875.00     4,275.00     9,150.00           18.09     88,207.13
Machine Setups setups       13,050.00                    1,050.00         950.00     2,000.00             6.53       6,851.25
Order Size MHs    502,027.00                    4,700.00     5,000.00     9,700.00           51.76 243,250.20
   680,635.00                  10,625.00 10,225.00 20,850.00           76.37 338,308.58
Estimated Production of Product A4           650.00
Cost per Unit           520.47